HUBBLE SPACE TELESCOPE DAILY REPORT #973 PERIOD COVERED: 0800 EST Daily Status Report as of 259/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed PC 4712 (WF/PC Cycle 2 Calibration: Measles Monitor) The Planetary Camera was used to make three Earth Calibration Flat images to monitor fo a form of contamination known as measles. The observations used filter F517N, both with and without filter F122M. The OSS monitored the observation and assessed the data quality. No problems were reported. All three images were reported to look okay. 1.2 Completed Two Sets of FGS 5183 (A Critical Test of the Galactic Escape Velocity at R(SUN), Continuation of 3856) The Fine Guidance Sensors were used to make astrometric observations of the high velocity star G-1625 to determine its galactic escape velocity. The FGSs was used in the POS mode to measure the target's trigonometric parallax and proper motion. Although only one STOP flag occurred, the FGS #3 scan pattern was seen for most of the of the targets were not what the co-investigator had expected to see. It was believed that there might be a problems with these data. 1.3 Completed Two Sets of WFC 4381 (HST Medium - Deep Survey - Cycle 3) The Wide Field Camera was used to make two set of medium-deep high latitude sky survey observations taken in parallel with the prime science observations. The observations were executed as scheduled, and no problems were reported. 1.4 Completed WFC 4770 (WF/PC Cycle 2 Calibration: Internal Monitor Proposal) The Wide Field Camera was used to make one bias and one dark calibration images. These data are required for routine calibration of the instrument and for data reduction purposes. Both of the observations were reported to be normal calibration images. 1.5 Completed PC 4308 (Snapshot Survey for Companions to Planetary - Nebula Nuclei - Cycle 3 High) The Planetary Camera was used to make an observation of the central region of planetary nebula VY2-2. It is possible that the central star in a double star. The observations was received and displayed in the OSS, the data was reported to be normal. 1.6 Completed GHRS/2 3472 (The Properties of Single Interstellar Clouds Cycle 2) The High Resolution Spectrograph was used to make Echelle spectral observations of the broad-lined, early-type star HD143018. Both the Echelle and grating G160M were used. All of the observations were executed, received and displayed. No problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 2 Successful : 2 Scheduled Reacquisitions : 3 Successful : 2 FGS #3 reacquisition failed to acquirer guide star in fine lock at 258/2018Z. It was caused by attitude errors as a result of type 59 slew. A loss of lock at 258/1858Z due to SIP being executed during fine lock on FGS#3. Reacquisitions attempted repeatedly at 258/1918Z ( about 10 times), but all failed. Subsequent acquisition was successful. 2.2 FHST Updates: Scheduled : 3 Successful : 3 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete WFC/PC 5026 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor) 3.2 Complete FOS/4396 (QSO Absorption System Snapshot Survey - Part 1 of 3: cycle 3 High) 3.3 Complete FGS 5183 (A Critical Test of the Galactic Escape Velocity at R(sun)) 3.4 Complete Four Sets of WFC 4381 (HST Medium - Deep Survey - Cycle 3) 3.5 Complete WFC/PC 4770 (WF/PC Cycle 2 Calibration: Internal Monitor Proposal) 3.6 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.7 Complete PC 5014 (WF/PC Cycle 3 calibration: PC Flats Narrow. 2) 3.8 Complete FOS/RD 4528 (SINS: The Supernova Intensive Study - SN1992) 3.9 Complete FOS 4276 (The abundance and Evolution of Carbon in the Spiral Galaxy NGC--2403) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-